The age of individual students have normal distribution with mean 18.6 and standard deviation 1.24. At Buting Senior High School, there 76 students from HUMSS A and HUMSS B with the same distribution as the entire age of grade 11.




What is the sample variance?

1
Expert's answer
2021-11-15T16:05:46-0500

Here, "\mu=18.6, \space \sigma=1.24, \space n=76"

Since the ages of 76 students from HUMSS A and HUMSS B has the same distribution as the entire age of grade 11 then, the sample standard deviation given by "\sigma_{\bar{x}}" is,

"\sigma _{\bar{x}}=\sigma/\sqrt n=1.24/\sqrt{76}= 0.1422378"

The sample variance is given as,

"\sigma^2_{\bar{x}}=(\sigma_{\bar{x}})^2=0.1422378^2= 0.02023158"

Therefore, the sample variance is 0.02023158.